# FPPL: An Efficient and Non-IID Robust Federated Continual Learning Framework

### v1

@@ Abstract
federated continual learning, prompt learning, prototype learning
@@ I Introduction
Federated Learning (FL) [29, 43, 18, 16] is considered as a specific distributed learning paradigm in which clients update their models by aggregating information from different client-side local models without uploading private data, thereby ensuring data privacy. Most existing FL methods assume that the data in each client is static without dynamically changing over time. However, in many real-world applications, the data of each client accumulates continuously, which would lead to the emergence of new classes or potential shifts in data distribution. Typical static FL methods become not applicable, the catastrophic forgetting issue, degrading performance on previous tasks, would significantly occur when data dynamically grows [20, 31], which could be exacerbated by the non-IID (non-independent and identically distributed) issue of FL [48, 24].
@@
Therefore, it becomes natural to consider the Federated Continual Learning (FCL) [44, 7]. FCL aims to establish federated learning methods that can adapt to new data streams and mitigate catastrophic forgetting [31] on previous tasks. However, prevalent methods often resort to rehearsal [7, 30], a common technique in continual learning [31, 41, 49], where data from prior tasks is utilized for subsequent training to mitigate catastrophic forgetting. This practice, integrating old task data into the current task, raises privacy concerns or adds storage and computational overhead in FCL. For instance, GLFC [7] locally store old task data for rehearsal, posing privacy risks and introducing additional burdens on clients. Moreover, FedCIL [30] uses generative replay to produce synthetic data, avoiding raw data maintenance but still incurring additional storage and computational costs due to the use of generative models. Rehearsal-based methods heavily rely on incorporating old task data into subsequent training, potentially leading to privacy issues or increased storage and training pressure on the local model. As a consequence, these methods lack scalability when faced with an increasing number of new tasks. Instead, the most efficient approach is to update the model by training solely on new data, but such a rehearsal-free method may result in catastrophic forgetting.
@@
Fortunately, the utilization of pre-trained models [9] has emerged as a promising solution to mitigate catastrophic forgetting without the need for rehearsal[40, 39, 32]. FCL methods utilizing pre-trained transformers [36, 6, 8] have gained prominence [27, 1]. The use of pre-trained transformers eliminates the need for rehearsal, providing an efficient alternative for mitigating catastrophic forgetting. For instance, recent work [1] proposes a rehearsal-free FCL approach based on pre-trained Vision Transformer [8], demonstrating a communication-efficient transmission strategy. Regrettably, methods based on pre-trained transformers have not sufficiently addressed the critical challenge of non-IID issue in federated learning [48, 24], lacking developed methodologies to mitigate resulting performance degradation effectively. In federated learning, the non-IID nature introduces data heterogeneity, leading to significant bias in models trained on different clients, and aggregating these biased models directly may result in a suboptimal global model [24, 46]. Similar to conventional federated learning, FCL encounters performance degradation with increasing non-IID degrees, as outlined in [47]. Additional knowledge sharing among clients is necessary to mitigate the impact of non-IID.
@@
To address the challenge of mitigating catastrophic forgetting without rehearsal, we draw inspiration from prompt-based continual learning methods [40, 39, 32], which utilize lightweight prompts [22, 25, 17] to instruct pre-trained transformers. The lightweight nature of prompts makes them well-suited for federated learning scenarios, where they can effectively reduce communication cost of clients, yielding promising performance without rehearsal. To tackle the issue of non-IID, lightweight prototypes [34], defined as the mean of extracted features from data, have been employed as efficient and privacy-preserving information carriers in federated learning [35, 15]. Building on the insights from [35, 15], we propose using prototypes as a superior alternative to rehearsal for addressing performance degradation caused by non-IID issue.
@@
In this work, an efficient and non-IID robust federated continual learning framework, called Federated Prototype-Augmented Prompt Learning (FPPL), is proposed. The proposed FPPL leverages both prompt tuning and prototypes to tackle the challenges of non-IID and catastrophic forgetting. Specifically, task-specific prompts are integrated with a fusion function to fully leverage acquired knowledge, thereby mitigating catastrophic forgetting. To address performance degradation caused by both non-IID data and catastrophic forgetting, a classifier debiasing mechanism based on local prototypes is designed. These local prototypes are aggregated into global prototypes at the server and returned to clients, achieving unified representations across clients and further addressing the non-IID issue. The main contributions can be summarized as:
@@ II-A Federated Continual Learning
The study of FCL has gained substantial traction in recent years. FedWeIT [44] utilizes decomposed global federated parameters and sparse task-specific parameters to minimize interference between incompatible tasks and allow positive knowledge transfer across clients. GLFC [7] designs class-aware and class-semantic loss for FCL, utilizing a proxy server to reconstruct local data at the server. Approaches [30, 47] leverage generative models to generate synthetic data for knowledge distillation [13] to alleviate catastrophic forgetting.
With the development of pre-trained models [9], FCL methods based on pre-trained transformers [36, 6, 8] have achieve notable performance with fewer communication cost. FedET [27] incorporates enhancer modules to address catastrophic forgetting issue while reducing communication overhead. Fed-CPrompt [1], based on prompting techniques, facilitates communication-efficient FCL without the need for rehearsal. While FCL methods leveraging pre-trained transformers have demonstrated notable performance with reduced communication cost, there exists a gap in the exploration of their performance across varying non-IID degrees.
@@ II-B Continual Learning
Continual learning, alternatively known as incremental learning or lifelong learning [4, 28, 38, 50], is a paradigm involving models acquiring knowledge from sequential tasks. Throughout the learning process of new tasks, models are either restricted from accessing or can only partially access data pertaining to prior tasks. The primary goal of continual learning is to mitigate the phenomenon of forgetting old tasks while concurrently acquiring knowledge on new ones. Continual learning can be broadly categorized into three classes: regularization-based methods, rehearsal-based methods, and model expansion-based methods.
@@
Recent advancements in continual learning include methods that leverage knowledge from pre-trained transformers. L2P [40] introduced a key-query matching mechanism for selecting prompts from a prompt pool to instruct the knowledge from the pre-trained model for learning new tasks. Building upon L2P, DualPrompt[39] enhanced performance by incorporating both general prompts and expert prompts. CODA-Prompt [32] utilizing an attention mechanism to construct input-conditioned prompts. While these prompt-based methods demonstrate applicability in continual learning, they are constrained by their reliance on centralized dataset specifics, limiting their applicability to FCL challenges.
@@ III Preliminaries
Furthermore, by expanding the setting of class-incremental learning to federated learning scenario, the $t$-th task’s training data $\mathcal{D}^{t}$ is distributed randomly among $K$ clients as $\mathcal{D}^{t}=\{{\mathcal{D}_{k}^{t}}\}_{k=1}^{K}$. It is assumed that the data heterogeneity across clients are distribution-based label skew [24, 46], and for any clients $k\neq k^{\prime}$, it holds that $\mathcal{D}_{k}^{t}\cap\mathcal{D}_{k^{\prime}}^{t}=\emptyset$. As for FCL, due to the limited computational and storage resources of clients and privacy guarantee, we focus on rehearsal-free continual learning setting [33].
@@ IV-A Local Prompt Tuning
During the learning process, the backbone parameters $\Phi$ of ViT remain frozen, while only the task-specific prompt $p_{t}$, coslinear $\psi$, and classifier parameters $\theta$ are properly tuned. These parameters are collectively expressed as $\bm{w}=\left\{\psi,p_{t},\theta\right\}$. At the first task ($t=1$), the task-specific prompt $p_{1}$ is uniformly initialized and then optimized. After completing a task ($t\leftarrow t+1$), the newly obtained task-specific prompt is initialized as the previous prompt ($p_{t+1}\leftarrow p_{t}$) to obtain acquired knowledge. To mitigate catastrophic forgetting, the prompts corresponding to previously completed tasks $\{p_{i}\}_{i=1}^{t-1}$ will be frozen. At task $t$, we will gather $P=\{p_{i}\}_{i=1}^{t}$, where $p_{t}$ denoting the $t$-th task-specific prompt. Since the newly obtained prompt is based on previous prompts, we utilize a tunable coslinear to derive weights for task-specific prompts. These weights are further integrated through a fusion function to fully exploit accumulated knowledge.
@@
The query is further processed by a cosine similarity-based linear layer $g:D\rightarrow\mathbb{R}^{t}$, mapping the query to a $t$-dimensional output based on cosine similarity, i.e., $g_{\psi}(q_{\mathbf{x}})=\frac{q_{\mathbf{x}}\psi}{\|q_{\mathbf{x}}\|_{2}\|\psi\|_{2}}\in\mathbb{R}^{t}$. The weights for each task-specific prompt are computed using the softmax function, and an element-wise multiplication operation is applied to derive the following fused prompt:
@@
where $\sigma(\cdot)$ denotes the softmax function and $\odot$ denotes the element-wise multiplication operation. The complete feature extraction function is denoted as $c_{\mathbf{x}}=f_{\Phi}\left(\left[\mathbf{x},\bm{p}\right]\right)$, where $\left[\mathbf{x},\bm{p}\right]$ is the combination of input and fused prompt, thus the final output logits is denoted as $o_{\mathbf{x}}=h_{\theta}(c_{\mathbf{x}})$ for classification. For convenience, the query, extracted feature and final output logits for input $\mathbf{x}$ is denoted as $q_{\mathbf{x}}$, $c_{\mathbf{x}}$ and $o_{\mathbf{x}}$ respectively.
@@ IV-B Global Prototype Aggregation
where $c_{\mathbf{x}}$ is the extracted feature of input $\mathbf{x}$ and $\mathcal{D}_{k}^{y}$ denotes the subset of private dataset of client $k$ with label $y$. After local training procedure in current round, the obtained model will extract local prototypes to form local prototype set $\bm{C}_{k}^{t}=\{\mathcal{C}_{k}^{y}\}_{y\in\mathcal{Y}_{k}^{t}}$ and $\mathcal{Y}_{k}^{t}$ represents the available classes of client $k$ at task $t$. These prototype sets are synchronously transferred to the server. Further, these locally uploaded prototype sets can be merged at the server with the prototype pool as $\bm{C}_{g}=\bm{C}_{g}\cup\{\bm{C}_{k}^{t}\}_{k=1}^{K}$, after the completion of current task $t$.
@@
It is worth emphasizing that the debias operation on the server side does not cause excessive computational and storage pressure, because of the lightweight of the prototype collection and parameters. The efficiency of this debias operation for performance degradation alleviation can be demonstrated in the following experiment section.
@@
where $\mathcal{Y}^{t}$ represents the global available classes at task $t$ and $c_{\mathbf{x}}$ is the extracted feature of input $\mathbf{x}$. $\textrm{sim}(\cdot,\cdot)$ denotes the cosine similarity function following [3, 23]. $\tau$ denotes the temperature that controls the tolerance of difference between extracted feature and corresponding global prototype. It is worth noting that the size of the returned global prototype is consistent with the local prototype generated by each client, so that no additional transmission burden is incurred.
The cross-entropy loss resulting from the classification of each client is denoted as
@@
where $o_{\mathbf{x}}$ is the final output logits of input $\mathbf{x}$ and $y$ denotes the label of input $\mathbf{x}$. To summarize the modeling process above, the overall loss function for each client is
@@ V-A Experimental Setup
To measure the degree of forgetting, the average forgetting is calculated as, $\bar{F}=\frac{1}{T-1}\sum_{i=1}^{T-1}\max_{t\in\{1,\cdots,T-1\}}(a_{i,t}-a_{i,T})$. Here, $\max_{t\in\{1,\cdots,T-1\}}(a_{i,t}-a_{i,T})$ mesures the gap between the accuracy of task $i$ at task $T$ and the optimal accuracy achievable for task $i$, capturing the forgetting of previous tasks after the whole learning process. The primary metric is average accuracy, with average forgetting serving as a supplementary metric; achieving low average forgetting alongside low average accuracy suggests poor adaptability to new tasks, which is undesirable in FCL.
@@
2:     Initialize parameters $\bm{w}=\{\psi,p,\theta\}$, global prototype set $\bm{G}=\{\}$ and prototype pool $\bm{C}_{g}=\{\}$;
@@
- SLCA [45], an advanced continual learning method using slow learner and classifier alignment.
- GLFC [7], a rehearsal-based federated continual learning method, stores selected samples on the client side and reconstructs local data on the server to identify the best old global model, which then supports local relation distillation.
TABLE I: Results on 20-task ImageNet-R and 20-task CUB-200. $\bar{A}$ gives the accuracy averaged over tasks, $A_{T}$ gives the accuracy of all tasks after the last incremental stage, $\bar{F}$ gives the average forgetting, and $COMM$ gives the communication cost per round. We report results over 3 trials.
@@
Recent findings by SLCA [45] suggest that the performance of full fine-tuning methods can be enhanced by utilizing a lower learning rate for the representation layer of the pre-trained model and a slightly higher rate for the classifier, referred to as the slow learner approach. Consequently, we enhance EWC, LwF and GLFC 111 The original paper [7] utilizes ResNet18 [10] as the backbone. For fairness, we replace it with the ViT-B/16 model pre-trained on ImageNet, aligning with other methods [40, 39, 32]. with slow learner approach advocated by SLCA.
To emphasis, we will implement these continual learning methods in the FCL scenario, while naming them as FedEWC, FedLwF, FedL2P, FedDualP, FedCODA-P and FedSLCA respectively.
Implementation Details. Consistent with prior works [40, 39, 32], we employ a ViT-B/16 model pre-trained on ImageNet as the common backbone for all baseline methods. Adam [19] optimizer is used for all methods with the same batch size of 128. Prompt-based methods uniformly use a constant learning rate of $lr=1e^{-3}$. Conversely, full fine-tuning methods use a learning rate of $lr_{rps}=1e^{-4}$ for the representation layer and $lr_{cls}=1e^{-2}$ for the classifier, in accordance with the slow learner proposed by SLCA [45]. We set clients $K=10$, overall communication rounds $R_{all}=200$ and local update epochs $E=5$, the communication rounds of each task is calculated by $R={\textstyle{\frac{R_{all}}{T}}}$ following  [7, 47]. For the FPPL, server update epochs $E_{server}=5$ and temperature $\tau=0.2$ empirically.

### v3

@@ Abstract
federated continual learning, prompt tuning, prototype learning
@@ I Introduction
Federated Learning (FL) [29, 43, 18, 16] is considered as a specific distributed learning paradigm in which clients update their models by aggregating information from different client-side local models without uploading private data, thereby ensuring data privacy. Most existing FL methods assume that the data of clients is static without dynamically changing over time. However, in real-world applications, the data of clients accumulates continuously, which would lead to the emergence of new classes or potential shifts in data distribution. Typical static FL methods become not applicable, the catastrophic forgetting issue, degrading performance on previous tasks, would significantly occur when data dynamically grows [20, 31], which could be exacerbated by the non-IID (non-independent and identically distributed) issue of FL [48, 24].
Therefore, it becomes natural to consider the Federated Continual Learning (FCL) [44, 7]. FCL aims to establish federated learning methods that can adapt to new data streams and mitigate catastrophic forgetting [31] on previous tasks.
@@
Prevalent FCL methods typically rely on rehearsal [7, 30], a common technique in continual learning [31, 41, 49], where data from prior tasks is utilized for subsequent training to mitigate catastrophic forgetting. This practice, integrating previous task data into the current task, raises privacy concerns or adds storage and computational overhead in FCL. For instance, GLFC [7] locally store previous task data for rehearsal, posing privacy risks and introducing additional burdens on clients. Moreover, FedCIL [30] uses generative replay to produce synthetic data, avoiding raw data maintenance but still incurring additional storage and computational cost due to the use of generative models. As a consequence, rehearsal-based methods lack scalability when faced with an increasing number of new tasks. Instead, the most efficient approach is to update the model by training solely on new data, but such a rehearsal-free method may result in catastrophic forgetting.
@@
Fortunately, leveraging pre-trained transformers [36, 6, 8] has proven to be a promising approach for mitigating catastrophic forgetting without the need for rehearsal. For instance, Fed-CPrompt [1] proposes a rehearsal-free FCL approach based on pre-trained Vision Transformer [8]. Regrettably, methods based on pre-trained models have not sufficiently addressed the critical challenge of non-IID issue in federated learning [48, 24], lacking developed methods to mitigate resulting performance degradation effectively. In federated learning, the non-IID nature introduces data heterogeneity, leading to significant bias in models trained on different clients, and aggregating these biased models directly may result in a suboptimal global model [24, 46]. Similar to conventional federated learning, FCL encounters performance degradation with increasing non-IID degrees, as outlined in [47]. Additional knowledge sharing among clients is necessary to mitigate the impact of non-IID.
@@
To tackle the issue of non-IID, lightweight prototypes [34], defined as the mean of extracted features from data, have been employed as efficient and privacy-preserving information carriers in federated learning [35, 15]. Building on the insights from [35, 15], we propose using prototypes as a superior alternative to rehearsal for addressing performance degradation caused by both catastrophic forgetting and non-IID issues. To reduce the communication cost associated with transmitting pre-trained transformers, we draw inspiration from prompt-based continual learning methods [40, 39, 32], which utilize lightweight prompts [22, 25, 17] to instruct models.
@@
In this work, an efficient and non-IID robust federated continual learning framework, called Federated Prototype-Augmented Prompt Learning (FPPL), is proposed. The proposed FPPL leverages both prompt tuning and prototypes to tackle the challenges of non-IID and catastrophic forgetting. Specifically, task-specific prompts are integrated with a fusion function to fully leverage acquired knowledge, thereby mitigating catastrophic forgetting. To address performance degradation caused by both non-IID data and catastrophic forgetting, a classifier debiasing mechanism based on local prototypes is designed. These local prototypes are aggregated into global prototypes on the server and returned to clients, achieving unified representations across clients through contrastive learning and further addressing the non-IID issue. The main contributions can be summarized as:
@@ II-A Federated Continual Learning
The study of FCL has gained substantial traction in recent years. FedWeIT [44] utilizes decomposed global federated parameters and sparse task-specific parameters to minimize interference between incompatible tasks and allow positive knowledge transfer across clients. GLFC [7] designs class-aware and class-semantic loss for FCL, and utilizes a proxy server to reconstruct local data. Approaches [30, 47] leverage generative models to generate synthetic data for knowledge distillation [13] to alleviate catastrophic forgetting.
With the development of pre-trained models [9], FCL methods based on pre-trained transformers [36, 6, 8] have achieve notable performance. FedET [27] incorporates enhancer modules to address catastrophic forgetting challenge while reducing communication overhead. Fed-CPrompt [1], based on prompts, facilitates communication-efficient FCL without the need for rehearsal.
While FCL methods using pre-trained transformers have demonstrated notable performance, their effectiveness across varying non-IID degrees remains underexplored.
@@ II-B Continual Learning
Continual learning, alternatively known as incremental learning or lifelong learning [4, 28, 38, 50], is a paradigm involving models acquiring knowledge from sequential tasks. Throughout the learning process of new tasks, models are either restricted from accessing or can only partially access data pertaining to prior tasks. The primary goal of continual learning is to mitigate the phenomenon of forgetting previous tasks while concurrently acquiring knowledge on new ones. Continual learning can be broadly categorized into three classes: regularization-based methods, rehearsal-based methods, and model expansion-based methods.
@@
Recent advancements in continual learning include methods that leverage knowledge from pre-trained transformers. L2P [40] introduced a key-query matching mechanism for selecting prompts from a prompt pool to instruct the knowledge from the pre-trained model for learning new tasks. Building upon L2P, DualPrompt[39] enhanced performance by incorporating both general prompts and expert prompts. CODA-Prompt [32] utilizing an attention mechanism to construct input-conditioned prompts.
@@ III Preliminaries
Furthermore, by expanding the setting of class-incremental learning to federated learning scenario, the $t$-th task’s training data $\mathcal{D}^{t}$ is distributed randomly among $K$ clients as $\mathcal{D}^{t}=\{{\mathcal{D}_{k}^{t}}\}_{k=1}^{K}$. It is assumed that the data heterogeneity across clients are distribution-based label skew [24, 46], and for any clients $k\neq k^{\prime}$, it holds that $\mathcal{D}_{k}^{t}\cap\mathcal{D}_{k^{\prime}}^{t}=\emptyset$.
@@ IV-A Local Prompt Tuning
During the learning process, the backbone parameters $\Phi$ of ViT remain frozen, while only the task-specific prompt $p_{t}$, coslinear parameters $\psi$, and classifier parameters $\theta$ are properly tuned. These parameters are collectively expressed as $\bm{w}=\left\{\psi,p_{t},\theta\right\}$. At the first task ($t=1$), the task-specific prompt $p_{1}$ is uniformly initialized and then optimized. After completing a task ($t\leftarrow t+1$), the newly obtained task-specific prompt is initialized as the previous prompt ($p_{t+1}\leftarrow p_{t}$) to obtain acquired knowledge.
To mitigate catastrophic forgetting, the prompts corresponding to previously completed tasks $\{p_{i}\}_{i=1}^{t-1}$ will be frozen following [39]. At task $t$, we will gather $P=\{p_{i}\}_{i=1}^{t}$, where $p_{t}$ denoting the $t$-th task-specific prompt. Since the newly obtained prompt is based on previous prompts, we utilize a tunable coslinear to derive weights for task-specific prompts. These weights are further integrated through a fusion function to fully exploit accumulated knowledge.
@@
The query is further processed by the coslinear layer $g_{\psi}$, mapping the query to a $t$-dimensional output based on cosine similarity, i.e., $g_{\psi}(q_{\mathbf{x}})=\frac{q_{\mathbf{x}}\psi}{\|q_{\mathbf{x}}\|_{2}\|\psi\|_{2}}\in\mathbb{R}^{t}$. The weights for each task-specific prompt are computed using the softmax function, and an element-wise multiplication operation is applied to derive the following fused prompt:
@@
where $\sigma(\cdot)$ denotes the softmax function and $\odot$ denotes the element-wise multiplication operation.
The complete feature extraction function is then denoted as $c_{\mathbf{x}}=f_{\Phi}\left(\left[\mathbf{x},\bm{p}\right]\right)$, in which $\left[\mathbf{x},\bm{p}\right]$ is the combination of input and fused prompt. Thus the final output logits is denoted as $o_{\mathbf{x}}=h_{\theta}(c_{\mathbf{x}})$ for classification.
For brevity, the query, extracted feature, and final output logits of input $\mathbf{x}$ are denoted as $q_{\mathbf{x}}$, $c_{\mathbf{x}}$, and $o_{\mathbf{x}}$, respectively.
@@ IV-B Global Prototype Aggregation
where $c_{\mathbf{x}}$ is the extracted feature of input $\mathbf{x}$ and $\mathcal{D}_{k}^{y}$ denotes the subset of private dataset of client $k$ with label $y$.
After local training procedure in current round, the obtained model will extract local prototypes to form local prototype set $\bm{C}_{k}^{t}=\{\mathcal{C}_{k}^{y}\}_{y\in\mathcal{Y}_{k}^{t}}$ and $\mathcal{Y}_{k}^{t}$ represents the available classes of client $k$ at task $t$. These prototype sets are synchronously transferred to the server. Further, these locally uploaded prototype sets can be merged on the server with the prototype pool as $\bm{C}_{g}=\bm{C}_{g}\cup\{\bm{C}_{k}^{t}\}_{k=1}^{K}$, after the completion of current task $t$, to retain knowledge from all tasks.
@@
It is worth emphasizing that the debias operation on the server side does not cause excessive computational and storage pressure, because of the lightweight of the prototype collection and parameters. The effectiveness of this debias operation for performance degradation alleviation can be demonstrated in the following experiment section.
@@
where $\mathcal{Y}^{t}$ represents the global available classes at task $t$ and $c_{\mathbf{x}}$ is the extracted feature of input $\mathbf{x}$. The function $\textrm{sim}(\cdot,\cdot)$ denotes the cosine similarity function following [3, 23]. $\tau$ denotes the temperature that controls the tolerance of difference between extracted feature and corresponding global prototype.
It is worth emphasizing that the size of the returned global prototype aligns with the local prototype generated by each client, ensuring that no extra transmission burden is introduced during the process.
To summarize the training process on each client, the cross-entropy loss for classification is expressed as:
@@
where $o_{\mathbf{x}}$ is the final output logits of input $\mathbf{x}$ and $y$ denotes the label of input $\mathbf{x}$.
The overall loss function combines the unified representation loss with the classification loss, yielding:
@@ V-A Experimental Setup
To measure the degree of forgetting, the average forgetting is calculated as, $\bar{F}=\frac{1}{T-1}\sum_{i=1}^{T-1}\max_{t\in\{1,\cdots,T-1\}}(a_{i,t}-a_{i,T})$. Here, $\max_{t\in\{1,\cdots,T-1\}}(a_{i,t}-a_{i,T})$ mesures the gap between the accuracy of task $i$ at task $T$ and the optimal accuracy achievable for task $i$, capturing the forgetting of previous tasks after the whole learning process.
@@
2:     Initialize $\bm{w}=\{\psi,p,\theta\}$, global prototype set $\bm{G}=\{\}$, and prototype pool $\bm{C}_{g}=\{\}$;
@@
- SLCA [45], an advanced continual learning method using slow learner and classifier alignment;
@@
- GLFC [7], a rehearsal-based federated continual learning method, stores selected samples on the client side and reconstructs local data on the server side.
TABLE I: Results on 20-task ImageNet-R and 20-task CUB-200. $\bar{A}$ gives the accuracy averaged over tasks, $A_{T}$ gives the accuracy of all tasks after the last incremental stage, $\bar{F}$ gives the average forgetting, and $COMM$ gives the communication cost per round.
We report results over 3 trials.
@@
Recent findings by SLCA [45] suggest that the performance of full fine-tuning methods can be enhanced by utilizing a lower learning rate for the representation layer of the pre-trained model and a slightly higher rate for the classifier, referred to as the slow learner approach. Consequently, we enhance EWC, LwF and GLFC 111 The original paper [7] utilizes ResNet18 [10] as the backbone. For fairness, we replace it with the ViT-B/16 model pre-trained on ImageNet, aligning with other methods [40, 39, 32]. with the slow learner.
To emphasis, we will implement these continual learning methods in the FCL scenario, while naming them as FedEWC, FedLwF, FedL2P, FedDualP, FedCODA-P, and FedSLCA, respectively.
Implementation Details. Consistent with prior works [40, 39, 32], we employ a ViT-B/16 model pre-trained on ImageNet as the common backbone for all baseline methods. Adam [19] optimizer is used for all methods with the same batch size of 128. Prompt-based methods uniformly use a constant learning rate of $lr=1e^{-3}$. Conversely, full fine-tuning methods use a learning rate of $lr_{rps}=1e^{-4}$ for the representation layer and $lr_{cls}=1e^{-2}$ for the classifier, in accordance with the slow learner proposed by SLCA [45]. We set clients $K=10$, overall communication rounds $R_{all}=200$ and local update epochs $E=5$, the communication rounds of each task is calculated by $R={\textstyle{\frac{R_{all}}{T}}}$ following  [7, 47]. For the FPPL, server update is performed over $E_{server}=5$ and temperature is set to $\tau=0.2$ empirically.